# Probabilistic Consistency in Machine Learning and Its Connection to Uncertainty Quantification

### v1

@@ 3 Fundamental Mathematical Concepts
Following on the perspective outlined above, an analysis of ML can be split along two parallel tracks: independent constructions of $\Pr[C]$ and $\Pr[r|C]$. The former problem has been rigorously studied in a variety of contexts spanning diagnostics and epidemiology to machine learning, so that we do not consider it here. We refer readers to works on prevalence estimation [48, 49, 50] and the quantitation problem [22, 23, 24, 25, 26, 27, 28, 29, 30], for example. Here we focus on connecting classifiers $\hat{C}$ and distributions $\Pr[r|C]$ in the binary context. We remind the reader that while many of the definitions in this section have standard counterparts in the literature, we modify them so as to extract additional structure needed for later analysis. We indicate differences where appropriate.
@@ 3.1 Classifiers Revisited: A Set Theoretic Perspective
$D_{1}^{\star}(q)$ =$=\{r:q_{1}P_{1}(r)>q_{2}P_{2}(r)\}\cup B^{\star}_{1}(q),$ (12a)
$D_{2}^{\star}(q)$ =$=\{r:q_{1}P_{1}(r)<q_{2}P_{2}(r)\}\cup B^{\star}_{2}(q),$ (12b)
@@
Proof of Lemma 1 is tedious but straightforward; it relies on monotonicity and the fact that the $D_{j}(q)$ are an equivalence class. The key steps leverage the facts that: (i) once a point $r$ switches class as a function of $\alpha$, it cannot switch back; and (ii) for a fixed $r$, we are free to choose the class label at affine prevalence at which $\hat{C}$ is discontinuous. The following result is a trivial consequence of Lemma 1.
@@
is interesting for two reasons. First, $B^{\star}(q)$ is a level-set of the ratio $P_{2}(r)/P_{1}(r)$, and thus it stands to reason that the latter quantity is special in some way. Second, this equation suggests promoting the affine-prevalence from the role of a free parameter to that of a function of $r$, which yields the value of $q$ for which $r\in B^{\star}(q)$. In other words, it seems reasonable to reinterpret the statement $r\in B^{\star}(q)$ as a “set-theoretic equation” for an unknown $q$, which we could express informally by writing “$r\in B^{\star}(q;r)$” for a fixed value of $r$. The solution to this “equation” is given by some function ${{\mathscr{q}}}(r)$, which we define below.
To do so, we consider the following definitions.
@@
  • The appearance of $\Gamma_{j,k}$ in Def. 5 is non-trivial because $R_{j,k}$ is indeterminate outside of this set. In the binary case, we can simply overcome this by setting $\Gamma=\Gamma_{1,2}$. However, in the multiclass case the situation is more complicated if the supports $\Gamma_{j,k}$ are partially disjoint as $j$ and $k$ vary. In particular, it is not clear how to define $R_{j,k}$ on $\Gamma_{j^{\prime},k^{\prime}}/\Gamma_{j,k}$ if $\Gamma_{j,k}\neq\Gamma_{j^{\prime},k^{\prime}}$, since the density ratio takes the indeterminate form $/0$ 0 on this set. Addressing this issue is a key task in Sec. 4.
  • For similar reasons, boundedness of the $P_{j}(r)$ is henceforth a critical assumption of our analysis. Because we do not assume continuity of the underlying densities, it is difficult, if not impossible, to resolve indeterminate forms such as $\infty$/$\infty$ / ∞, which we would need to address if we did not assume boundedness.
@@
Let $U^{\star}(q)\cong\hat{C}^{\star}(r,q)$. Fix a value of $r$. Assume first that $P_{j}(r)>0$ for $j=1,2$. By Prop. 1, there exists a $\tilde{q}$ for which $\tilde{q}_{1}\in(0,1)$ such that $r\in B^{\star}(\tilde{q})$. By Lemma 1 and Def. 4, we know that for $q_{1}>\tilde{q}_{1}$, the point $r\in D^{\star}_{1}(q)$, and hence $\hat{C}^{\star}(r,q)=1$. For $q_{1}<\tilde{q}_{1}$, one finds $\hat{C}^{\star}(r,q)=2$. Since $\hat{C}^{\star}(r,q)$ is binary and monotone in $q$ for fixed $r$, there can only be one point of discontinuity, which must be $\tilde{q}$. Equality of the supremum and infimum also follows by monotonicity.
@@ 4.1 Constructing Bayes Classifiers from Pairwise Counterparts
$D_{j}^{\star}(q)$ =$=\{r:q_{j}P_{j}(r)>q_{k}P_{k}(r),\forall k,k\neq j\}\cup B_{j}^{\star}(q)$
=$=\{r:q_{j}/q_{k}>R_{j,k}(r),\forall k,k\neq j\}\cup B_{j}^{\star}(q),$ (19)
@@
To make use of this observation, it is helpful to “project” a multiclass classifier onto its pairwise counterparts via the following definition.
@@
In contrast with Def. 1, the definition of a pairwise classifier does not require that $j<k$. However, note that there are only $K(K-1)/2$ unique pairwise classifiers for $\hat{C}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$, and similarly for the $R_{j,k}$ and ${{\mathscr{q}}}_{j,k}$, and we can construct all such functions by only considering those for which $j<k$.
In practice, it is generally easier to construct a pairwise classifier than its multiclass counterpart. Thus it is reasonable to assume that one always has access to the pairwise Bayes classifier $\hat{C}^{\star}_{j,k}$, from which Prop. 2 yields the relative density ratios $R_{j,k}$. We now show that this information is sufficient for both constructing the multiclass Bayes classifier and estimating the inherent uncertainty associated with any population.
@@
We construct a Bayes classifier $\hat{C}^{\star}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$ in the multiclass setting using only pairwise training. Assume that the support $\Gamma_{j}$ of each $P_{j}(r)$ is the full measurement space, i.e. $\Gamma=\Gamma_{j}$. Let $\{\hat{C}_{j,k}^{\star}\}$ be the $K(K-1)/2$ unique pairwise Bayes classifiers. By Proposition 2, for any point $r$, we can determine ${{\mathscr{q}}}_{j,k}(r)$ in terms of the value of $\alpha$ at which the $\hat{C}_{j,k}^{\star}(r,\alpha)$ is discontinuous.
Consider next a test population $\Omega(\chi)$ with test prevalence $\chi$. Given the pairwise prevalence functions, we can construct the pointwise accuracy of an arbitrary classifier $\hat{C}:\Gamma\to{\mathbb{K}}$ as
$Z(r;\chi,\hat{C})$ $=\sum_{j=1}^{K}\frac{\chi_{j}P_{j}(r)}{Q(r;\chi)}\mathbb{I}\left(\hat{C}(r)=j\right)$
@@
$\hat{C}(r)=\operatorname*{argmax}_{j}\left[\frac{\chi_{j}P_{j}(r)}{Q(r;\chi)}\right]=\operatorname*{argmax}_{j}\left[\frac{\chi_{j}}{\sum_{k}\chi_{k}{{\mathscr{q}}}_{j,k}(r)/[1-{{\mathscr{q}}}_{j,k}(r)]}\right].$ (21)
Moreover, observe that the resulting $\hat{Z}(r;\chi,\hat{C})=1-u(r)$ yields the inherent uncertainty, and thus $\hat{C}$ is the Bayes classifier for test prevalence $\chi$. That is, $\hat{C}(r)=\hat{C}^{\star}(r,\chi)$. Q.E.F.
A key, and perhaps surprising, prediction of Construction 1 is that when the supports of the $P_{j}(r)$ fully overlap, a Bayes classifier never need be trained on all classes simultaneously, only pairs thereof. This is sometimes referred to as an all-pairs or a one-versus-one training paradigm [59, 60, 61, 62, 63]; see also Ref. [64] and the concept of reduction. This approach also stands in contrast to other popular methods such as one-versus-all [65, 66]; see also Refs. [67, 68]. Construction 1 was foreshadowed in Refs. [20, 33] but does not appear to have been proved rigorously. Here it is critical to note that a straightforward extension of binary class-switching to the multiclass setting relies on the rather restrictive assumption that the $\Gamma_{j}=\Gamma$, which is not reasonable in many settings.
@@
To accomplish this, fix $r$ and let ${\mathscr{C}}_{m,n}^{\star}$ be arbitrary extensions of the pairwise classifiers onto $\Gamma$, with $\tilde{{\mathscr{q}}}_{m,n}$ the corresponding extensions of the prevalence function. Clearly there must be at least one value of $m$ for which $r\in\Gamma_{m}$. Assume first that $r\notin\Gamma_{n}$ for $n\neq m$. Note that ${{\mathscr{q}}}_{m,n}(r)$ and ${{\mathscr{q}}}_{n,m}(r)$ are defined on $\Gamma_{m,n}=\Gamma_{m}\cup\Gamma_{n}$, so that $\tilde{{\mathscr{q}}}_{m,n}(r)={{\mathscr{q}}}_{m,n}(r)$ and $\tilde{{\mathscr{q}}}_{n,m}(r)={{\mathscr{q}}}_{n,m}(r)$ on this set. Extend Eq. (21) to $\Gamma$ according to
@@ 4.2 Determining When a Classifier is (not) Bayes Optimal
Up to this point, we have analyzed the properties of Bayes classifiers, tacitly assuming that these are the outputs of a training paradigm. In particular, we have assumed existence of the densities $P_{j}(r)$ and used these to study properties of $\hat{C}^{\star}$. But we may also consider whether a set of arbitrary pairwise classifiers implies the existence of density functions. In other words, we ask whether it is possible to know a priori (i.e. without access to a training population) that a classifier $\hat{C}$ is actually Bayes optimal across all of $\Gamma\times{{\rm X}^{K}}$. Moreover, we seek to deduce properties of the conditional densities from $\hat{C}$.
The motivation for this question arises from the fact that classifiers are often constructed from imperfect training processes involving finite data and/or incomplete optimization. Thus, it is useful to have techniques to assess (i) whether a training population is “induced” by a classifier, and (ii) if that population is “close” to the available training data. Moreover, in applications such as image classification, it is not clear that the $P_{j}$ could even be represented in an intelligible way, so that when $\hat{C}^{\star}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$ exists, it is in fact the representation of those densities. Equally important, it is useful to have necessary and sufficient conditions that guarantee a set of pairwise classifiers can be used to construct a multiclass Bayes classifier in the spirit of Construction 2. Such conditions could inform numerical optimization algorithms in ML and address unresolved issues of probabilistic consistency discussed in Refs. [32].
As of yet, we are unaware of a complete solution to this problem, and herein we only present partial results. In particular, it is straightforward to prove that monotone binary classifiers with “normalizable” density ratios are always Bayes classifiers for which we can extract density functions. We also argue that the general problem can be restated in terms of constrained optimization. The issue amounts to the fact that the $P_{j}(r)$ are required to be non-negative, which is non-trivial to guarantee in the multiclass case. In the following sections, we therefore limit ourselves to (i) deriving additional properties of Bayes classifiers, (ii) showing that these are necessary conditions for an arbitrary $\hat{C}$ to induce densities $P_{j}$, and (iii) examining special cases.
@@ 4.2.1 Law of Total Probability for Classifiers
$R_{1,3}(r)$ =1=1,= 1 , (28c)
for $\epsilon\ll 1$ clearly fail to satisfy Eq. (26). We can interpret the first two ratios as stating that $r$ is “not in class 2” and “not in class 3,” but the last density ratio asserts that classes 1 and 3 have equal probability. Thus it appears that $r$ does not belong in any of the classes, which is nominally a contradiction. In other words, Eq. (26) ensures that the classifier yields a valid interpretation: if $K-1$ classes have low probability, the last must have high probability.
@@
Let $\hat{R}_{j,k}:\Gamma\to[0,\infty]$ for $j,k\in{\mathbb{K}}$. Then the $\hat{R}_{j,k}$ can be expressed in terms of at most $K$ non-negative, bounded functions via Eq. (15) if and only if the $\hat{R}_{j,k}(r)$ are self-consistent.
@@
In the event that one or more of the $\hat{R}_{j,k}$ takes the values of 0 or $\infty$, self-consistency implies that there exists a subset $W\subset{\mathbb{K}}$ such that the density ratios are positive and bounded. By the previous argument, we may clearly define the $\hat{R}_{j,k}$ as ratios of the form $\hat{R}_{j,k}=P_{k}/P_{j}$ for $j,k\in W$. Next consider $m\in V$. The choice $P_{m}=0$ agrees with the assignment $\hat{R}_{j,m}=P_{m}/P_{j}=0$ and $\hat{R}_{m,j}=P_{j}/P_{m}=P_{j}/0=\infty$. $\qed$
@@
Let $\hat{C}:\Gamma\times{{\rm X}^{K}}$ be a self-consistent classifier, and let $\hat{R}_{j,k}$ and $\hat{{\mathscr{q}}}_{j,k}$ be the induced density ratios. Then the induced prevalence functions satisfy Eq. (25). Moreover, if there exists a Bayes classifier $\hat{C}^{\star}$ having density ratios $\hat{R}_{j,k}$, then $\hat{C}$ is that classifier. In particular, the inherent uncertainty $u(r)$ is given by Eq. (20) with ${{\mathscr{q}}}_{i,j}(r)=\hat{{\mathscr{q}}}_{i,j}(r)$.
@@
To show that $\hat{C}$ is a Bayes classifier, let $\hat{C}^{\star}:\hat{\Gamma}\times{{\rm X}^{K}}$ be any Bayes classifier constructed from densities $P_{j}(r)$ whose supports are $\hat{\Gamma}_{j}$ and having density ratios $R_{i,j}(r)=\hat{R}_{i,j}(r)$, where $R_{i,j}:\hat{\Gamma}_{i,j}\to[0,\infty]$. Clearly this classifier is self-consistent and has the same sets $W(r)$ and $V(r)$ as $\hat{C}$. Hence $\hat{C}$ is the Bayes classifier. $\qed$

### v2

@@ 3 Fundamental Mathematical Concepts
Following on the perspective outlined above, an analysis of ML can be split along two parallel tracks: independent constructions of $\Pr[C]$ and $\Pr[r|C]$. The former problem has been rigorously studied in a variety of contexts spanning diagnostics and epidemiology to machine learning, so that we do not consider it here. We refer readers to works on prevalence estimation [48, 49, 50] and the quantitation problem [22, 23, 24, 25, 26, 27, 28, 29, 30], for example. Here we focus on establishing an equivalence between classifiers $\hat{C}$ and distributions $\Pr[r|C]$ in the binary context. We remind the reader that while many of the definitions in this section have standard counterparts in the literature, we modify them so as to extract additional structure needed for later analysis. We indicate differences where appropriate.
Before proceeding, it is also useful to clarify that our main task can be further divided into two questions: (i) given a Bayes-optimal classifier, can we extract information about the true underlying $\Pr[r|C]$; and (ii) given an arbitrary classifier $\hat{C}$, is there a natural way to associate with it distributions $\Pr[r|C]$? This section focuses the first question in a binary setting; Section 4 addresses both questions in a multiclass setting.
@@ 3.1 Classifiers Revisited: A Set Theoretic Perspective
$D_{1}^{\star}(q)$ $=\{r:q_{1}P_{1}(r)>q_{2}P_{2}(r)\}\cup B^{\star}_{1}(q),$ (12a)
$D_{2}^{\star}(q)$ $=\{r:q_{1}P_{1}(r)<q_{2}P_{2}(r)\}\cup B^{\star}_{2}(q),$ (12b)
@@
Proof of Lemma 1 is tedious but straightforward; it relies on monotonicity and the fact that the $D_{j}(q)$ are an equivalence class. The key steps leverage the facts that: (i) once a point $r$ switches class as a function of $\alpha$, it cannot switch back; and (ii) for a fixed $r$, we are free to choose the class label at the affine prevalence at which $\hat{C}$ is discontinuous. The following result is a trivial consequence of Lemma 1.
@@
is interesting for two reasons. First, $B^{\star}(q)$ is a level-set of the ratio $P_{2}(r)/P_{1}(r)$, and thus it stands to reason that the latter quantity is special in some way. Second, this equation suggests promoting the affine-prevalence from the role of a free parameter to that of a function of $r$, which yields the value of $q$ for which $r\in B^{\star}(q)$. In other words, it seems reasonable to reinterpret the statement $r\in B^{\star}(q)$ as a “set-theoretic equation” for an unknown $q$, which we could express informally by writing $r\in B^{\star}(q;r)$” for a fixed value of $r$. The solution to this “equation” is given by some function ${{\mathscr{q}}}(r)$, which we define below. In particular, consider the following.
@@
  • The appearance of $\Gamma_{j,k}$ in Def. 5 is non-trivial because $R_{j,k}$ is indeterminate outside of this set. In the binary case, we can simply overcome this by setting $\Gamma=\Gamma_{1,2}$. However, in the multiclass case the situation is more complicated if the supports $\Gamma_{j,k}$ are partially disjoint as $j$ and $k$ vary. In particular, it is not clear how to define $R_{j,k}$ on $\Gamma_{j^{\prime},k^{\prime}}/\Gamma_{j,k}$ if $\Gamma_{j,k}\neq\Gamma_{j^{\prime},k^{\prime}}$, since the density ratio takes the indeterminate form $0/0$ on this set. Addressing this issue is a key task in Sec. 4.
  • For similar reasons, boundedness of the $P_{j}(r)$ is henceforth a critical assumption of our analysis. Because we do not assume continuity of the underlying densities, it is difficult, if not impossible, to resolve indeterminate forms such as $\infty/\infty$, which we would need to address if we did not assume boundedness.
@@
[…]
@@ 4 Class-Switching in the Multiclass Setting
The class-switching representation of the density ratios demonstrates how to extract probabilistic information from binary Bayes classifiers. While the $R_{j,k}(r)$ do not fully specify all of the distributional information in the $P_{j}(r)$, in the binary case, they are sufficient constructing $u(r)$.
In this context, the purposes of the present section are twofold. First, we wish to generalize the results of the previous section to a multiclass setting. In particular, we show that given a multiclass Bayes classifier in the sense of Eq. (6), we can still extract sufficient probabilistic information to compute aleatoric uncertainty. A surprising outcome of this exercise is the identification of two additional properties, self-consistency and normalizability, that all Bayes classifiers satisfy. This leads to the second purpose of the present section, which is to study the extent to which these properties, along with monotonicity, are sufficient for assigning a meaningful probabilistic interpretation to arbitrary classifiers; see also the discussion at the beginning of Sec. 3.
@@ 4.1 Constructing Bayes Classifiers from Pairwise Counterparts
$D_{j}^{\star}(q)$ $=\{r:q_{j}P_{j}(r)>q_{k}P_{k}(r),\forall k,k\neq j\}\cup B_{j}^{\star}(q)$
$=\{r:q_{j}/q_{k}>R_{j,k}(r),\forall k,k\neq j\}\cup B_{j}^{\star}(q),$ (19)
@@
These observations suggest that we consider the restriction of $\hat{C}^{\star}$ to the boundary $\partial{{\rm X}^{K}}$ of the probability simplex as follows.
@@
In contrast with Def. 1, the definition of a pairwise classifier does not require that $j<k$. However, note that there are only K(K1)/2K(K-1)/2 unique pairwise classifiers for $\hat{C}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$, and similarly for the $R_{j,k}$ and ${{\mathscr{q}}}_{j,k}$, and we can construct all such functions by only considering those for which $j<k$.
In practice, it is generally easier to construct a pairwise classifier than its multiclass counterpart. Thus it is reasonable to assume that one always has access to the $\hat{C}^{\star}_{j,k}$, from which Prop. 2 yields the relative density ratios $R_{j,k}$. We now show that this information is sufficient for both constructing the multiclass Bayes classifier and estimating the inherent uncertainty associated with any population.
@@
We construct a Bayes classifier $\hat{C}^{\star}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$ in the multiclass setting using only pairwise training. Assume that the support $\Gamma_{j}$ of each $P_{j}(r)$ is the full measurement space, i.e. $\Gamma=\Gamma_{j}$. Let $\{\hat{C}_{j,k}^{\star}\}$ be the K(K1)/2K(K-1)/2 unique pairwise Bayes classifiers. By Proposition 2, for any point $r$, we can determine ${{\mathscr{q}}}_{j,k}(r)$ in terms of the value of $\alpha$ at which the $\hat{C}_{j,k}^{\star}(r,\alpha)$ is discontinuous.
Consider next a test population $\Omega(\chi)$ with test prevalence $\chi$ in the interior of ${\rm X}^{K}$}. Given the pairwise prevalence functions, we can construct the pointwise accuracy of an arbitrary classifier $\hat{C}:\Gamma\to{\mathbb{K}}$ as
[…]
@@
$\hat{C}(r)=\operatorname*{argmax}_{j}\left[\frac{\chi_{j}P_{j}(r)}{Q(r;\chi)}\right]=\operatorname*{argmax}_{j}\left[\frac{\chi_{j}}{\sum_{k}\chi_{k}{{\mathscr{q}}}_{j,k}(r)/[1-{{\mathscr{q}}}_{j,k}(r)]}\right],$ (21)
where we may assign $\hat{C}(r)$ to any class attaining the maximum in the event that more than one $j$ solves Eq. (21). Moreover, observe that the resulting $\hat{Z}(r;\chi,\hat{C})=1-u(r)$ yields the inherent uncertainty, and thus $\hat{C}$ is the Bayes classifier for test prevalence $\chi$. That is, $\hat{C}(r)=\hat{C}^{\star}(r,\chi)$. Q.E.F.
A key, and perhaps surprising, outcome of Construction 1 is that when the supports of the $P_{j}(r)$ fully overlap, a Bayes classifier never need be trained on all classes simultaneously, only pairs thereof. This is sometimes referred to as an all-pairs or a one-versus-one training paradigm [59, 60, 61, 62, 63]; see also Ref. [64] and the concept of reduction. This approach also stands in contrast to other popular methods such as one-versus-all [65, 66]; see also Refs. [67, 68]. Construction 1 was foreshadowed in Refs. [20, 33] but does not appear to have been proved rigorously. Here it is critical to note that a straightforward extension of binary class-switching to the multiclass setting relies on the rather restrictive assumption that the $\Gamma_{j}=\Gamma$, which is not reasonable in many settings.
@@
To accomplish this, fix $r$ and let ${\mathscr{C}}_{m,n}^{\star}$ be arbitrary extensions of the pairwise classifiers onto $\Gamma$, with $\tilde{{\mathscr{q}}}_{m,n}$ the corresponding extensions of the prevalence function. As before, let $\chi$ be in the interior of ${\rm X}^{K}$}. Clearly there must be at least one value of $m$ for which $r\in\Gamma_{m}$. Assume first that $r\notin\Gamma_{n}$ for some $n\neq m$. Note that ${{\mathscr{q}}}_{m,n}(r)$ and ${{\mathscr{q}}}_{n,m}(r)$ are defined on $\Gamma_{m,n}=\Gamma_{m}\cup\Gamma_{n}$, so that $\tilde{{\mathscr{q}}}_{m,n}(r)={{\mathscr{q}}}_{m,n}(r)$ and $\tilde{{\mathscr{q}}}_{n,m}(r)={{\mathscr{q}}}_{n,m}(r)$ on this set. Extend Eq. (21) to $\Gamma$ according to
@@ 4.2 Determining When a Classifier is (not) Bayes Optimal
Up to this point, we have analyzed the properties of Bayes classifiers, tacitly assuming that these are the outputs of a training paradigm. In particular, we have assumed existence of the densities $P_{j}(r)$ and used these to study properties of $\hat{C}^{\star}$. But we may also consider whether a set of arbitrary pairwise classifiers implies the existence of density functions. To accomplish this, we ask whether it is possible to know a priori (i.e. without access to a training population) that a classifier $\hat{C}$ is actually Bayes optimal across all of $\Gamma\times{{\rm X}^{K}}$. Moreover, we seek to deduce properties of the conditional densities from $\hat{C}$.
The motivation for this question arises from the fact that classifiers are often constructed from imperfect training processes involving finite data and/or incomplete optimization. Thus, it is useful to have techniques to assess (i) whether a training population is “induced” by a classifier, and (ii) if that population is “close” to the available training data. Lacking the first property, it is not clear that meaningful uncertainty estimates can be assigned to a classifier; see Sec. 2.2. Moreover, in applications such as image classification, it is often unclear how to express $P_{j}$ in an intelligible way, so that when $\hat{C}^{\star}:\Gamma\times{{\rm X}^{K}}\to{\mathbb{K}}$ exists, it is in fact the representation of those densities. Equally important, it is useful to have necessary and sufficient conditions that guarantee a set of pairwise classifiers can be used to construct a multiclass Bayes classifier in the spirit of Construction 2. Such conditions could inform numerical optimization algorithms in ML and address unresolved issues of probabilistic consistency discussed in Refs. [32].
As of yet, we are unaware of a complete solution to this problem, and herein we only present partial results. In particular, it is straightforward to prove that monotone binary classifiers with “normalizable” density ratios are always Bayes classifiers for which we can extract density functions. We also argue that the general problem can be restated in terms of constrained optimization. The issue amounts to the fact that the $P_{j}(r)$ are required to be both normalizable and non-negative, which is non-trivial to guarantee in the multiclass case. In the following sections, we therefore limit ourselves to (i) deriving additional properties of Bayes classifiers, (ii) showing that these are necessary conditions for an arbitrary $\hat{C}$ to induce densities $P_{j}$, and (iii) examining special cases.
@@ 4.2.1 Law of Total Probability for Classifiers
$R_{1,3}(r)$ $=1,$ (28c)
for $\epsilon\ll 1$ clearly fail to satisfy Eq. (26). We can interpret the first two ratios as stating that $r$ is “not in class 2” and “not in class 3,” but the last density ratio asserts that classes 1 and 3 have equal probability. Thus it appears that $r$ does not belong in any of the classes, which is nominally a contradiction. In other words, Eq. (26) ensures that the classifier yields a valid interpretation: if K1K-1 classes have low probability, the last must have high probability.
@@
Let $\hat{R}_{j,k}:\Gamma\to[0,\infty]$ for $j,k\in{\mathbb{K}}$. Then excluding indeterminate forms, the $\hat{R}_{j,k}$ can be expressed in terms of at most $K$ non-negative, bounded functions via Eq. (15) if and only if the $\hat{R}_{j,k}(r)$ are self-consistent.
@@
In the event that one or more of the $\hat{R}_{j,k}$ takes the values of 0 or $\infty$, self-consistency implies that there exists a subset $W\subset{\mathbb{K}}$ such that the density ratios are positive and bounded. By the previous argument, we may clearly define the $\hat{R}_{j,k}$ as ratios of the form $\hat{R}_{j,k}=P_{k}/P_{j}$ for $j,k\in W$. Next consider $m\in V$. The choice $P_{m}=0$ agrees with the assignment $\hat{R}_{j,m}=P_{m}/P_{j}=0$ and $\hat{R}_{m,j}=P_{j}/P_{m}=P_{j}/0=\infty$, with the remaining density ratios indeterminate (i.e. taking any value). $\qed$
@@
Let $\hat{C}:\Gamma\times{{\rm X}^{K}}$ be a self-consistent classifier, and let $\hat{R}_{j,k}$ be the induced density ratios. Then the induced prevalence functions $\hat{{\mathscr{q}}}_{j,k}$ satisfy Eq. (25). Moreover, if there exists a Bayes classifier $\hat{C}^{\star}$ having density ratios $\hat{R}_{j,k}$, then $\hat{C}$ is that classifier. In particular, the inherent uncertainty $u(r)$ is given by Eq. (20) with ${{\mathscr{q}}}_{i,j}(r)=\hat{{\mathscr{q}}}_{i,j}(r)$.
@@
Finally, let $\hat{C}^{\star}:\hat{\Gamma}\times{{\rm X}^{K}}$ be any Bayes classifier constructed from densities $P_{j}(r)$ whose supports are $\hat{\Gamma}_{j}$ and having density ratios $R_{i,j}(r)=\hat{R}_{i,j}(r)$, where $R_{i,j}:\hat{\Gamma}_{i,j}\to[0,\infty]$. Clearly this classifier is self-consistent and has the same sets $W(r)$ and $V(r)$ as $\hat{C}$. Hence $\hat{C}$ is the Bayes classifier. $\qed$